OpenStudy (anonymous):

Use the percent equation to answer: What number is 92% of 115? Can someone show me how to solve with percent equations?

OpenStudy (anonymous):

0.92*115

OpenStudy (katrinakaif):

"What number" = n Is always stands for "equals" 92% becomes .92 Of = x or multiplication What number is 92% of 115? n = .92 x 115

OpenStudy (katrinakaif):

And solve

OpenStudy (unklerhaukus):

\[90% \times 115 \] \[=90/100 \times 115\] \[=103.5\]

OpenStudy (anonymous):